OpenStudy (anonymous):

true or false The solution of the equation 2x = 49 is approximately 5.61. i don't even remember where to start….

OpenStudy (uri):

We start by solving for x; what does x equal if x = 49/2?

OpenStudy (anonymous):

24.5

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

\[2x = 49\]\[ \frac{ 2x }{ 2 } =\frac{ 49 }{ 2 }\]\[x=24.5\]

OpenStudy (uri):

@caitlynxoxo__98 that's correct, so then the answer is false xD